# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ Abstract.
Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization. Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conduct the Systematic Mapping Study (SMS) on the literature published between January 2013 and March 2023. Eighty-six studies were finally selected and the key results are: (1) Four leading themes on cloud computing and network systems research were identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Seventy-four patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four frameworks were introduced from the analysis of identified challenges and solutions: security, automation, deployment, and monitoring challenge-solution frameworks. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions.
Containerization, Multi-Cloud Environment, Challenges, Solutions, Frameworks, Systematic Mapping Study
journal: TOSEM
@@ 1. Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b). Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014). Containers allow the packaging of an application along with its required dependencies, which makes it easy to move the application across different environments with little or no modification (Docker, 2021). On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013). Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
Containerization in multi-cloud is illustrated in Figure 1. This figure provides an overview of applications deployed within various cloud configurations including public, private, and hybrid models. Each cloud hosts multiple applications encapsulated in containers, and completes different types of tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization offers. A container platform layer, which could be exemplified by systems like Kubernetes or Docker, manages and orchestrates these containers across the different cloud environments. This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
@@
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022). These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may involve ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018). Similarly, In the system implementation phase, challenges may include dealing with potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
@@
Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and the grey literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive solutions.
@@
To address this knowledge gap, we established the objective of this research is to systematically identify and categorize various facets of container utilization in multi-cloud environment. These facets encompass the roles of containers, their implementation strategies, architectural patterns, quality attributes, as well as the motivations of the customers. Additionally, we aspire to explore the challenges and solutions related to automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis.
@@
The empirical findings and frameworks of this Systematic Mapping Study (SMS) offer significant potential benefits for developers of container-based applications. By exploring a diverse range of implementation strategies, architectural patterns, and quality attributes, practitioners can gain valuable insights to enhance their practices. Moreover, a detailed examination of challenges and solutions related to automation, deployment, monitoring, and security further provides practitioners with the knowledge to address problems effectively and improve dedicated quality attributes. Additionally, these findings and frameworks offer valuable support to practitioners in comprehending various patterns, strategies, and tactics specific to containerized applications in multi-cloud environment. The proposed frameworks can act as a facilitator of knowledge sharing, training, and ensure a shared understanding among practitioners regarding the complexities and solutions within containerized applications in multi-cloud environment. Ultimately, this SMS serves to empower software developers and practitioners, enabling them to optimize their approaches and elevate the quality and efficiency of containerized applications in multi-cloud environment.
Our SMS makes the following key contributions:
@@
  • We developed four frameworks to address challenges and solutions in security, automation, deployment, and monitoring for containerization in multi-cloud environment.
@@
  • We systematically identified and classified an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
@@
  • We systematically identified and classified quality attributes and tactics for containerized applications in multi-cloud environment.
@@
  • We identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
@@
  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access all collected data, replicate, and validate our SMS. This dataset includes detailed hierarchies of the developed frameworks, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
@@ 2. Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007). Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process.
@@ 2.1. Research Questions
To conduct this SMS, we formulated six Research Questions (RQs), as outlined in Table 1, aligning with the objective of our SMS. The table provides a structured view of the RQs, their corresponding rationale, and categorization.
@@
RQ1
What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment?
To investigate the role of containers and explore implementation strategies in multi-cloud environment, with the objective of enhancing understanding and facilitating efficient deployment and management of applications across multiple cloud environments.
RQ2
What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment?
To explore and document the various architectural patterns and strategies that are utilized in implementing container-based applications in multi-cloud environment, thus providing a roadmap for efficient and scalable application development and deployment.
RQ3
What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment?
To examine and analyze the quality attributes required for container-based applications in multi-cloud environment and the corresponding tactics employed in their implementation, enabling the identification of best practices and guidelines for developing robust and reliable applications in such environments.
@@
RQ4
What are the challenges related to security, automation, deployment, and monitoring in container-based applications in multi-cloud environment?
To comprehensively evaluate and gain insights into the challenges surrounding security, automation, deployment, and monitoring concerning container-based applications in multi-cloud environment, this question aims to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing these applications in such a dynamic setting.
RQ5
How to address security, automation, deployment, and monitoring in container-based applications within multi-cloud environment?
This question seeks to investigate solutions aimed at tackling the challenges of security, automation, deployment, and monitoring within container-based applications operating in multi-cloud environment. By providing actionable guidance and recommendations, the goal is to elevate the overall management and performance of these applications in the context of multi-cloud environment.
@@
RQ6
What tools and frameworks are used to implement container-based multi-cloud applications?
To investigate and identify the specific tools and frameworks utilized in the implementation of container-based multi-cloud applications, providing an overview of the technological landscape and aiding developers and organizations in making informed decisions when selecting appropriate tools for their application deployment and management.
@@ 2.2.1. Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and March 2023 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environments over the past decade. We followed the steps mentioned below to select the relevant studies.
@@
Language
English
Non-English
Study Type
Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers
Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
Study Focus
Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
Study Duration
Studies published between January 2013 and March 2023
Studies published before January 2013 or after March 2023
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 5,676 studies from eight databases.Following initial retrieval, we did not apply Cohen’s Kappa as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. Initially, we removed any duplicated studies obtained from different databases (such as IEEE Xplore and Scopus) by sorting them in ascending order. This resulted in the elimination of several hundred duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objectives, as detailed in Table 3. To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1968) was performed on a random sample of 40 studies, as illustrated in Formula 1:
@@
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of 0.531. This value reflects a moderate level of agreement between R1 and R2. Any disagreements were systematically resolved through discussions among all authors, leading to a consensus. Following this process, the number of studies was reduced to 809 for further analysis.
@@
  • Step 3: Abstract-based Screening: During this step, we conducted a thorough review of the abstracts of the collected studies to determine their relevance to our research topic. Two researchers (R1 and R2) independently examined each abstract, assigning a status of ”relevant,” ”irrelevant,” or ”doubtful” with respect to our research goals. To reduce subjective bias and to quantify the consensus between the researchers, we conducted a Cohen’s Kappa analysis on a subset of 30 randomly selected studies. The outcomes of this assessment are depicted in Table 4.b, which show substantial agreement between R1 and R2, with 25 studies being unanimously considered relevant and only 4 studies disagree (3 with one researcher favoring relevance and the other not, and 1 where both disagreed) The Kappa value, calculated based on the observed agreements and disagreements was approximately K= 0.268 which suggests a fair agreement between the two researchers. Consequently, upon the compilation of step, we got 303 studies for more detailed scrutiny.
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 303 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly chose 20 studies (see Table 4.c) for a Cohen’s Kappa analysis to measure the agreement between two researchers. The resulting Kappa value was K=0.318, which represents a fair level of agreement. This measure confirms that our selection criteria were applied consistently during the full-text screening studies.
@@
(a) Title and Keyword Screening on 40 Randomly Selected Studies
@@
R2 Yes 35 2
R2 No 1 2
Total 36 4
(b) Abstract-based Screening on 30 Randomly Selected Studies
@@
R2 Yes 25 3
R2 No 1 1
Total 26 4
(c) Full-Text Screening on 20 Randomly Selected Studies

### v2

@@ Abstract.
Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization. Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conducted the Systematic Mapping Study (SMS) on the literature published between January 2013 and July 2024. One hundred twenty one studies were selected and the key results are: (1) Four leading themes on containerization in multi-cloud environment are identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Ninety-eight patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four catalogs consisting of challenges and solutions related to security, automation, deployment, and monitoring were introduced. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions for containerization applications in multi-cloud environment.
Containerization, Multi-Cloud Environment, Systematic Mapping Study, Cloud Computing
journal: TOSEMccs: Software and its engineering Software development techniques
@@ 1. Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b). Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014). Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification (Docker, 2021). On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013). Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost-efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
Containerization in multi-cloud is illustrated in Figure 1. This figure provides an overview of applications deployed within various cloud configurations, including public, private, and hybrid models. Each cloud hosts multiple applications encapsulated in containers, which can support different types of applications, such as those in healthcare or finance, that require flexibility in deployment and scalability across cloud providers. These containers complete various tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization provides. A container platform layer, which could be exemplified by systems like Kubernetes or Docker, manages and orchestrates these containers across the different cloud environments. For instance, a healthcare application can use a private cloud for processing sensitive data for compliance and a public cloud for data analysis to achieve cost-efficiency and scalability. This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
@@
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022). These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018). Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
@@
Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design (Ahmad et al., 2025). The project brings together a consortium of four companies (Bittium [^1] https://www.bittium.com/, M-Files [^2] https://www.m-files.com/, Solita [^3] https://www.solita.fi/en/, and Vaadin [^4] https://vaadin.com/), each with distinct requirements for containerization, and is further supported by industry leaders Nokia [^5] https://www.nokia.com/ and TietoEVRY [^6] https://www.tietoevry.com/. This study fulfills an industry demand for optimized container strategies, addressing critical gaps in deployment consistency, security, and scalability that are necessary for practical applications in multi-cloud contexts. Despite the growing interest in containerization for multi-cloud environment, there remains a significant gap in systematically organized knowledge to address these challenges comprehensively. Our study not only fulfills an immediate industry demand but also aims to provide a structured view of containerization practices, addressing the needs of both practitioners and researchers in navigating this fragmented knowledge space. Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and gray literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive, secure solutions.
@@
To bridge this knowledge gap and support the practical demands of the QLEAP project consortium, our study systematically identifies and categorizes the various facets of container utilization in multi-cloud environment. These facets include container roles, implementation strategies, architectural patterns, and quality attributes. In addition, we explore the challenges and solutions associated with automation, deployment, monitoring, and security, providing practitioners and researchers with actionable insights to tackle these critical issues effectively. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis. We adopted the following Goal-Question-Metric (GQM) framework (Caldiera and Rombach, 1994):
  • Goal: To enhance the understanding and efficiency of deploying and managing container-based applications in multi-cloud environment.
  • Questions: To explore specific aspects such as container roles, implementation strategies, architectural patterns, quality attributes, challenges, and solutions.
  • Metrics: To evaluate the effectiveness of our findings based on the identified roles, strategies, patterns, attributes, challenges, solutions, and tools.
@@
The empirical findings and challenge-solution catalogs from this Systematic Mapping Study (SMS) provide significant benefits for developers of container-based applications. By examining implementation strategies, architectural patterns, and quality attributes, practitioners gain insights to improve practices. Detailed analyses of challenges and solutions in automation, deployment, monitoring, and security equip practitioners to address issues and enhance quality attributes. These challenge-solution catalogs facilitate knowledge sharing, training, and a shared understanding of containerized applications in multi-cloud environments, empowering developers to optimize their approaches and improve efficiency and quality. In response to the real-world needs identified in the QLEAP project and the broader research gap, our SMS presents the following key contributions:
@@
  • We systematically identified and categorized challenges and their solutions related to security, automation, deployment, and monitoring for containerization in a multi-cloud environment.
@@
  • We systematically identified and categorized an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
@@
  • We systematically identified and categorized quality attributes and tactics for containerized applications in multi-cloud environment.
@@
  • We systematically identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
@@
  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS. This dataset includes detailed hierarchies of the developed catalogs, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
@@ 2. Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007). Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process. Reliability and consistency are critical aspects of any empirical study. In this study, we followed established statistical approaches, including inter-rater agreement (IRA) computation, to ensure the reliability of our data coding and extraction processes. Additionally, we applied systematic quality assessment techniques to evaluate the studies included in our SMS.
@@ 2.1. Research Questions
To conduct this SMS, we formulated six Research Questions (RQs) based on the GQM approach presented in the Introduction section (see Section 1) aligns with the objective of our SMS, as outlined in Table 1. The table provides a structured view of the RQs with their corresponding rationale and categorization.
@@
RQ1 What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment? To investigate the role of containers and explore implementation strategies in multi-cloud environment, with the objective of enhancing understanding and facilitating efficient deployment and management of applications across multiple cloud environments.
RQ2 What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment? To explore and document the various architectural patterns and strategies that are utilized in implementing container-based applications in multi-cloud environment, thus providing a roadmap for efficient and scalable application development and deployment.
RQ3 What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment? To examine and analyze the quality attributes required for container-based applications in multi-cloud environment and the corresponding tactics employed in their implementation, enabling the identification of best practices and guidelines for developing robust and reliable applications in such environments.
@@
RQ4 Challenges in Container-Based Applications in Multi-Cloud Environment RQ4.1:What are the challenges related to security in container-based applications in multi-cloud environment? RQ4.2: What are the challenges related to automation in container-based applications in a multi-cloud environment? RQ4.3: What are the challenges related to deployment in container-based applications in a multi-cloud environment? RQ4.4: What are the challenges related to monitoring in container-based applications in a multi-cloud environment? To comprehensively evaluate and gain insights into the challenges surrounding security (RQ4.1), automation (RQ4.2), deployment (RQ4.3), and monitoring (RQ4.4) concerning container-based applications in multi-cloud environment. These sub-RQs aim to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing container-based applications in a multi-cloud environment.
RQ5 Solutions to Address Challenges in Container-Based Applications in Multi-Cloud Environment RQ5.1:How to address the challenges related to security in container-based applications within a multi-cloud environment? RQ5.2: How to address the challenges related to automation in container-based applications within a multi-cloud environment? RQ5.3: How to address the challenges related to deployment in container-based applications within a multi-cloud environment? RQ5.4: How to address the challenges related to monitoring in container-based applications within a multi-cloud environment? These sub-RQs seek to investigate solutions aimed at tackling the challenges of security (RQ5.1), automation (RQ5.2), deployment (RQ5.3), and monitoring (RQ5.4) within container-based applications operating in multi-cloud environment. By providing actionable guidance and recommendations, the goal is to elevate the overall management and performance of these applications in the context of multi-cloud environment.
@@
RQ6 What tools and frameworks are used to implement container-based multi-cloud applications? To investigate and identify the specific tools and frameworks utilized in the implementation of container-based multi-cloud applications, providing an overview of the technological landscape and aiding developers and organizations in making informed decisions when selecting appropriate tools for their application deployment and management.
These RQs are designed to address important gaps in the current understanding of container-based applications in multi-cloud environment. RQ1–RQ3 focus on understanding the roles of containers, implementation strategies, architectural patterns, and quality attributes. These questions aim to provide a comprehensive perspective, building on and organizing fragmented insights from prior studies. RQ4 and RQ5 focus on specific challenges and corresponding solutions related to security, automation, deployment, and monitoring in container-based multi-cloud environment. These questions aim to systematically address issues that are often only partially explored in existing works. Finally, RQ6 investigates tools and frameworks used for implementation, helping bridge the gap between academic research and practical applications. Together, these RQs provide a structured framework that enhances understanding and supports the development and management of container-based applications in multi-cloud environment, advancing the state-of-the-art in this domain
2.2. Search String Composition
Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework (Schardt et al., 2007) (Ralph et al., 2020) for constructing our search string. However, given the complexity and breadth of our topic, applying PICO strictly required an extensive array of terms and logical operators, exceeding the query limits of databases like ScienceDirect. Therefore, we developed the search string for this study based on authors’ knowledge and iterative trial searches. This approach allowed us to tailor the search strategy to the specific goal of our study, ensuring comprehensive coverage within practical constraints.
@@ 2.3.1. Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and July 2024 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environment over the past decade. We followed the steps mentioned below to select the relevant studies.
@@
Language English Non-English
Study Type Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
Study Focus Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment. Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
Study Duration Studies published between January 2013 and July 2024 Studies published before January 2013 or after July 2024
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 6,727 studies from eight databases. Following initial retrieval, we did not apply Cohen’s Kappa (Cohen, 1960) (Cohen, 1968) as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. Initially, we removed any duplicate studies obtained from different databases (e.g., IEEE Xplore, Scopus) by sorting them in ascending order. This resulted in the elimination of several thousand duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objective, as detailed in Table 3.
    To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1960) (Cohen, 1968) was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
@@
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 (Sim and Wright, 2005). Specifically, R1 and R2 agreed on 360 studies being relevant (“YES”) and 20 studies being irrelevant (“NO”). Disagreements occurred in 20 studies, which were systematically resolved through discussions among all authors, leading to a final consensus. Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency (Keele et al., 2007; Chandler et al., 2019). Following this process, the number of studies was reduced to 1,047 for further analysis.
@@
  • Step 3: Abstract-based Screening: During this step, we conducted a thorough review of the abstracts of the collected studies to determine their relevance to our research topic. Two researchers (R1 and R2) independently examined each abstract, assigning a status of “relevant”, “irrelevant”, or “doubtful” with respect to our research goals. To reduce subjective bias and to quantify the consensus between the researchers, we performed a Cohen’s Kappa analysis on a random subset of 105 studies (approximately 10% of the 1,047 studies remaining after the Title and Keyword Screening phase).
    The outcomes of this assessment are depicted in Table 4.b. The analysis showed substantial agreement between R1 and R2, with 88 studies being unanimously considered relevant and 5 studies unanimously considered irrelevant. Disagreements occurred in 12 cases (7 studies where one researcher marked “relevant” while the other marked “irrelevant”, and 5 studies initially classified as “doubtful”). The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement (Sim and Wright, 2005; Keele et al., 2007).
    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant”, “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation. Specifically:
    • Studies classified as “relevant” were mapped to “YES”.
    • Studies classified as “irrelevant” were mapped to “NO”.
    • Studies classified as “doubtful” were resolved through discussions between R1 and R2. If consensus was reached on “relevant”, the study was marked as “YES”; otherwise, it was marked as “NO”.
    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines (Keele et al., 2007; Chandler et al., 2019). After resolving all disagreements through systematic discussions among the authors, a final consensus was reached for the subset. Consequently, upon completing this step, the number of studies was reduced to 346 for more detailed scrutiny.
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 346 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly selected a subset of 35 studies (10% of the total 346 studies) for a Cohen’s Kappa analysis to measure the agreement between two researchers (R1 and R2). The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers (Sim and Wright, 2005). Specifically, R1 and R2 agreed on 30 studies as relevant (“YES”) and 2 studies as irrelevant (“NO”). Disagreements occurred in 3 studies (2 marked as “YES” by one researcher and “NO” by the other, and 1 marked as “NO” by both but flagged as uncertain). All disagreements were systematically resolved through discussions among the authors, leading to a final consensus. This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process (Chandler et al., 2019). At the end of this step, we obtained a total of 110 studies, which were deemed relevant for further analysis.
@@
(a) Title & Keyword Screening on 400 Randomly Selected Studies
@@
R2 Yes 360 20
R2 No 10 10
Total 370 30
(b) Abstract-based Screening on 105 Randomly Selected Studies
@@
R2 Yes 88 7
R2 No 5 5
Total 93 12
(c) Full-Text Screening on 35 Randomly Selected Studies